# HumanSplat: Generalizable Single-Image Human Gaussian Splatting with Structure Priors

### v1

@@ 4.3 Ablation Study
Latent Reconstruction Transformer. HumanSplat leverages the Stable Diffusion latent space for feature extraction. Experimental results demonstrate that utilizing the compact latent space for reconstructions leads to better rendering quality compared to using the pixel space, as detailed in Tab. 2(c) (a). We also find that patching in the time dimension for latent decoding by temporal VAE SVD:2023 adversely affects the quality, so we employ the vanilla VQ-VAE VQGAN:2023 decoder in this work.
[…]
Method Category THuman2.0 THuman2.0:2021 Twindom Twindom Rec. Time
[…]
PIFu PIFu:2019 18.093 0.911 0.137 - - - $\sim$30 Sec
LGM LGM:2024 20.013 0.893 0.116 19.840 0.851 0.292 $\sim$9.5 Sec
GTA GTA:2024 18.050 - - 17.669 0.741 0.418 $\sim$5.6 Min
SIFU SIFU:2024 22.025 0.921 0.084 19.714 0.832 0.312 $\sim$5.7 Min
SIFU SIFU:2024 22.102 0.923 0.079 - - - $\sim$5.7 Min
Magic123 magicanimate:2023 14.501 0.874 0.145 - - - $\sim$1 Hours
HumanSGD HumanSGD:2023 17.365 0.895 0.130 - - - $\sim$7 Min
TeCH TECH:2023 25.211 0.936 0.083 21.192 0.884 0.188 $\sim$4.5 Hours
HumanSplat (Ours) 24.033 0.918 0.055 23.346 0.913 0.125 $\sim$9.3 Sec
Figure 4: Qualitative comparison of ours against TeCH TECH:2023 , GTA GTA:2024 and LGM LGM:2024 on Thuman2.0 THuman2.0:2021 , Twindom Twindom and in-the-wild images. Our method achieves the highest quality. Note that TeCH achieves clearer results but fails to preserve the face ID.
Human Geometric Prior. To evaluate the robustness of HumanSplat to SMPL parameters, we employed various initialization strategies for the geometry. The results of these different methods are presented in Table 2(c) (b). In cases where the SMPL initialization involved extreme randomness, our model demonstrated robustness against SMPL estimation errors, as our approach of treating human priors as keys and values. Furthermore, the projection window $k_{win}$ ensures that parameters estimated from a single view detector PIXIE:2021 , as well as results from the multi-view version of SMPLify Pavlakos2019_smplifyx , remain within acceptable tolerances. The effectiveness of the human geometric prior is also demonstrated qualitatively in Figure 5 (a).
Window Size. As analyzed in Sec. 3.4.2, human priors are crucial for the robustness of the human reconstruction model. As demonstrated in Tab. 2(c) (c), the PSNR metric is significantly improved from 22.625db to 24.293db. When the window size $K_{win}$ is set to 1 (Project Only), the model struggles to handle loose clothing. We also explore how the performance of the model varies with different window sizes by setting $K_{win}=2$, in a latent space with a resolution of $64\times 64$.
@@
Semantics-guided Training Objectives. Thanks to the incorporation of semantic-guided hierarchical loss, HumanSplat can generate textures of higher quality for both the face and hand, while also promote the convergence of the training process, as depicted in Figure 5 (a) and (b). Furthermore, Figure 5 (c) demonstrates that HumanSplat further produces 3D consistent segmentation results, which have potential applications in areas such as editing and specific generation of individual parts of 3D humans.
@@
Figure 5: (a) Qualitative evaluation of human prior and hierarchical loss $\mathcal{L_{H}}$. PSNR/SSIM values are shown at the top of each image, presenting the effectiveness of the two strategies in the full pipeline. (b) LPIPS scores during training on the 2K2K validation set. (c) HumanSplat can provide 3D consistent segmentation results for potential applications.
@@
Table 2: Ablation study of several proposed designs on 2K2K  2k2k dataset.
@@
Temporal VAE SVD:2023 24.133 0.902 0.047
Ours rombach2022high 24.293 0.915 0.040
@@
Random 19.894 0.876 0.366
Muti-view 24.633 0.935 0.025
Ours PIXIE:2021 24.293 0.915 0.040
(c) Different project methods.
@@
Project Only 23.333 0.916 0.057
Ours($K_{win}=2$) 24.293 0.915 0.040
@@ 5 Conclusion
We present HumanSplat, a pioneering generalizable Human Gaussian Splatting network that derives 3D Gaussian Splatting properties from a single image. This model integrates generative diffusion and latent reconstruction transformer models with human structure priors, enhanced by a tailored semantic-aware hierarchical loss. These innovations achieve realistic and detailed reconstructions in a fast feed-forward manner without any optimization or fine-tuning, especially in the important focal areas such as the face and hands. Extensive experiments demonstrate that HumanSplat surpasses existing state-of-the-art methods in both quality and efficiency, including robust performance in handling challenging poses and loose clothing. This capability opens up a broad spectrum of potential applications.
Limitations and Future Works. Our methods have some limitations that can be addressed in future works: (a). While effective with most attire, the model may struggle with intricate garments and accessories. Future enhancements should incorporate 2D data and innovative training strategies, which is a promising direction for improving the handling of diverse and unusual clothing styles. (b). Although already efficient, increasing the computational speed by constraining Gaussian’s number or combining compact 3DGS with texture representation could further facilitate real-time applications, particularly on mobile devices. (c). Animating the reconstructed human models currently requires post-processing. Future work could introduce canonical space to reconstruct animatable avatars directly, streamlining the process and improving efficiency.
@@ Appendix A Implementation Details
A.1 Additional ovel View Synthesizer Details
The conditions of Difffusion model. SVD and SV3D SVD:2023 ; voleti2024sv3d replace CLIP text embeddings CLIP:2021 with the image embedding of the conditioning. Compared to text prompts, image prompts provide information more accurately and directly, and also save time that would otherwise be spent on text inversion or image caption  TECH:2023 ; SIFU:2024 .
Transfomer Block. The CLIP embedding matrix from the conditioning image is fed into the cross-attention layers of each transformer block, acting as the key and value, while the feature at that layer serves as the query. Moreover, the camera trajectory and the diffusion noise timestep are incorporated into the residual blocks.
@@ A.1 Additional ovel View Synthesizer Details
[…]
@@ A.2 Additional Latent Reconstruction Transformer Details
The network architecture of the proposed latent reconstruction transformer is shown in Fig. 6. The framework is composed of a latent encoder with intra- and inter-attention and a Gaussian parameter prediction module. It takes initial multi-view latent embeddings as input and performs projection-aware attention to aggregate human geometric information. From each output token, we decode the attributes of pixel-aligned Gaussians in the corresponding patch with a Conv $1\times 1$ layer.
@@
A.3 Additional Training Details
[…]
Table 3: Evaluation of different window sizes $K_{win}$ on the 2K2K evaluation dataset.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o Human Prior 22.635 (-1.658) 0.893 (-0.022) 0.182 (+0.142)
$K_{win}=1$ 23.333 (-0.960) 0.916 (-0.001) 0.057 (+0.017)
$K_{win}=3$ 24.383 (+0.089) 0.931 (+0.016) 0.041 (+0.001)
$K_{win}=4$ 24.013 (-0.280) 0.922 (+0.007) 0.048 (+0.008)
$K_{win}=2$ (Ours) 24.293 0.915 0.040
Appendix B More Results.
In Tab. 3, we present results for the 2K2K dataset across various window sizes. The results show that employing a window size of $k_{\text{win}}=1$ significantly enhances performance, improving the PSNR metric from 22.635 db to 23.333 db. Further refinement with a window size of $k_{\text{win}}=3$ continues to enhance the results, whereas using $k_{\text{win}}=4$ does not yield additional improvements. After considering both computational complexity and metrics, we selected a window size of $K_{win}=2$.
As for novel view and novel pose rendering, more results are shown in Fig. 7 and Fig. 8. For a single input image, we use AnimateAnyone AnimateAnyone:2023 for novel pose synthesis and HumanSplat for novel view synthesis. This functionality enables dynamic timestamps and real-time rendering of novel views, thereby enhancing immersive virtual exploration. Please also kindly refer to our supplementary video for more results.
@@ Appendix C Broader Impacts
Figure 8: Qualitative 4D Gaussian Splatting results on In-the-wild images, including novel view and pose rendering results.

### v2

@@ 4.3 Ablation Study
Latent Reconstruction Transformer. HumanSplat leverages the Stable Diffusion latent space for feature extraction. Experimental results demonstrate that utilizing the compact latent space for reconstructions leads to better rendering quality compared to using the pixel space, as detailed in Tab. 2(c) (a). We also find that patching in the time dimension for latent decoding by temporal VAE [95] adversely affects the quality, so we employ the vanilla VAE [94] decoder in this work.
Figure 5: Qualitative results showcasing reconstructions of humans in challenging poses, diverse identities, and varying camera viewpoints from in-the-wild images.
Figure 6: Qualitative Comparison of ours against HumanSGD [22] on in-the-wild images.
Human Geometric Prior. Tab. 2(c) (b) illustrates the significance of human geometric prior in our approach. By utilizing SMPL as the geometric prior for human representation, we assess the robustness of HumanSplat with respect to different SMPL parameters. We report both the estimated [2] and Ground Truth SMPL parameters [92] and compare the model that does not incorporate a human structure prior. These experiments provide insight into the benefit of incorporating a human structure prior, as significant performance degradation is observed without it. Furthermore, thanks to treat human priors as keys and values, HumanSplat exhibits robustness in testing against challenging poses with imperfect estimation, with only minor declines in metrics. The effectiveness of the human geometric prior is also demonstrated qualitatively in Fig. 7 (a).
Window Size. As analyzed in Sec. 3.4.2, human priors are crucial for the robustness of the human reconstruction model. As evidenced in Tab. 2(c) (c), when the window size $K_{\text{win}}$ is set to 1 (dubbed “Projection Only”), employing human priors significantly enhances performance, improving the PSNR metric from 22.635 to 23.333. We further explore setting window size to $K_{\text{win}}=2$ in a latent space with a resolution of $64\times 64$, which further improves the model’s PSNR metric from 23.333 to 24.294. This improvement is attributed to Humansplat with $K_{\text{win}}=2$ is capable of handling loose clothing and tolerating imprecise SMPL parameters. In Tab. 3, we present results for the 2K2K dataset across various window sizes. Further refinement with a window size of $K_{\text{win}}=3$ continues to enhance the results, whereas using $K_{\text{win}}=4$ does not yield additional improvements.
@@
Semantics-guided Training Objectives. Thanks to the incorporation of semantic-guided hierarchical losses, HumanSplat can generate higher-quality textures for both the face and hand, while also promoting the convergence of the training process, as depicted in Fig. 7 (a) and (b). Furthermore, Fig. 7 (c) demonstrates that HumanSplat further produces 3D coherent segmentation results, which have potential applications in editing and specific generation of individual parts of 3D humans.
@@
Figure 7: (a) Qualitative evaluation of human prior and hierarchical loss $\mathcal{L_{H}}$. PSNR/SSIM values are shown at the top of each image, presenting the effectiveness of the two strategies in the full pipeline. (b) LPIPS scores during training on the 2K2K validation set. (c) HumanSplat can provide 3D coherent segmentation results for potential applications.
@@
Table 2: Ablation study of several proposed designs on 2K2K [109] evaluation dataset.
@@
Temporal VAE [95] 24.133 0.902 0.047
Ours [94] 24.293 0.915 0.040
@@
GT SMPL 24.633 0.935 0.025
w/o SMPL 22.635 0.893 0.182
Ours [2] 24.293 0.915 0.040
(c) Different projection methods.
@@
Projection Only 23.333 0.916 0.057
Ours ($K_{\text{win}}=2$) 24.293 0.915 0.040
Table 3: Evaluation of different window sizes $K_{\text{win}}$ on the 2K2K [109] evaluation dataset.
Method PSNR $\uparrow$ SSIM $\uparrow$ LPIPS $\downarrow$
w/o Human Prior 22.635 (-1.658) 0.893 (-0.022) 0.182 (+0.142)
$K_{\text{win}}=1$ 23.333 (-0.960) 0.916 (-0.001) 0.057 (+0.017)
$K_{\text{win}}=3$ 24.383 (+0.089) 0.931 (+0.016) 0.041 (+0.001)
$K_{\text{win}}=4$ 24.013 (-0.280) 0.922 (+0.007) 0.048 (+0.008)
$K_{\text{win}}=2$ (Ours) 24.293 0.915 0.040
@@ 5 Conclusion
We present HumanSplat, a pioneering generalizable human reconstruction network that derives 3D Gaussian Splatting properties from a single image. This model integrates generative diffusion and latent reconstruction Transformer models with human structure priors, enhanced by a tailored semantic-aware hierarchical loss. These innovations achieve high-fidelity reconstruction results in a feed-forward manner without any optimization or fine-tuning, especially in the important focal areas such as the face and hands. Extensive experiments demonstrate that HumanSplat surpasses existing state-of-the-art methods in both quality and efficiency, including robust performance in handling challenging poses and loose clothing. This capability opens up a broad spectrum of potential applications.
Limitations and Future Works. Our method has some limitations that can be addressed in future works: (a). While effective with most attire, the model may struggle with intricate garments and accessories. Future enhancements should incorporate 2D data and innovative training strategies, which is a promising direction for improving the handling of diverse and unusual clothing styles. (b). Although already efficient, increasing the computational speed by constraining Gaussian’s number or combining compact 3DGS with texture representation could further facilitate real-time applications, particularly on mobile devices. (c). Animating the reconstructed human models currently requires post-processing. Future work could introduce canonical space to reconstruct animatable avatars directly, streamlining the process and improving efficiency.
@@ Appendix A Implementation Details
A.1 Additional Novel View Synthesizer Details
[…]
Transformer Block. The CLIP embeddings obtained from the conditioning image is fed into the cross-attention layers of each Transformer block, acting as keys and values, while the feature at that layer serves as the queries. Moreover, the camera trajectory and the diffusion noise timestep are incorporated into the residual blocks.
@@ A.1 Additional Novel View Synthesizer Details
3D Consistency Evaluation. We also conduct an evaluation of the 3D consistency of our novel-view synthesizer using the advanced local correspondence matching algorithm, MASt3R [122]. This approach involves one-to-one image matching between input views and their generated novel views, with the average number of matching correspondences serving as our metric. Our synthesizer exhibits a notable improvement in 3D consistency for human novel-view generation after fine-tuning with human datasets. Specifically, the original SV3D achieves an average of 723.25 matching points, while our novel-view synthesizer increases this number to 930.13. In comparison, the ground truth demonstrates a matching point number of 1106.33. This evaluation highlights that fine-tuning on human datasets significantly enhances multi-view consistency metrics.
@@ A.2 Additional Latent Reconstruction Transformer Details
Network Architecture. The network architecture of the proposed latent reconstruction Transformer is shown in Fig. 8. The framework is composed of a latent encoder with intra- and inter-attention and a Gaussian parameter prediction module. It takes initial multi-view latent embeddings as input and performs projection-aware attention to aggregate human geometric information. From each output token, we decode the attributes of pixel-aligned Gaussians in the corresponding patch with a Conv $1\times 1$ layer.
Hierarchical Loss. Our approach is efficient and can be utilized online without pre-computation. We calculate the visible face triangles given the mesh and camera parameters. Then each visible triangular face is assigned the corresponding semantic label [103]. Additionally, it provides labels with superior 3D consistency and avoids issues associated with training instability. First, the hierarchical loss is specifically focused on the head and hands, excluding hair and clothing, which can be further supervised using reconstruction loss $\mathcal{L}_{\text{Rec}}$. Second, instead of using segmentation to reconstruct and then concatenate different components, we design our loss function based on part segmentation. Thank to this design, HumanSplat is tolerant of inaccuracies in segmentation.
@@
Appendix B More Results
For practical applications demanding high input view reconstruction quality, which could potentially be directly derived from the input image. We improve the weights of the reconstruction loss for input views (referred to as reweighting loss), demonstrating that the current model with reweighting loss exhibits sufficient capacity to render input views with higher fidelity, as illustrated in Fig.9. We also conduct an ablation experiment for the reweighting loss function on the 2K2K Dataset [109]. There are slight declines in quantitative metrics, specifically, HumanSplat with reweighting loss achieves a numerical change of -0.141, -0.015, and +0.006 in PNSR, SSIM, and LPIPS. This suggests that reweighting loss is not “a bag of freebies” and HumanSplat without reweighting loss strikes a balance between the quality of novel views and the input views.
As for novel view and novel pose rendering, more results are shown in Fig. 10 and Fig. 11. For a single input image, we use Open-AnimateAnyone [123] for novel pose synthesis and HumanSplat for novel view synthesis. This functionality enables dynamic timestamps and real-time rendering of novel views, thereby enhancing immersive virtual exploration. Please also kindly refer to our supplementary video for more results.
@@ Appendix C Broader Impacts
Figure 9: Ablation Study on Reweighting Loss. Comparison of the original HumanSplat, trained with reweighting loss, and SIFU featuring texture.
Figure 11: Qualitative 4D Gaussian Splatting results on In-the-wild images, including novel view and pose rendering images. (Please zoom in for a detailed view)